# Diffusion Self-Weighted Guidance for Offline Reinforcement Learning

### v1

@@ D.5 Scaling of inference time in SWG
D.6 Computational resources
Our implementation is based on the jaxrl repository [23] using the JAX [3] and Flax [13] libraries. We also provide an implementation of our method in PyTorch [38]. All experiments were conducted on a 12GB NVIDIA RTX 3080 Ti GPU. Training takes approximately 0.5 hours for value function learning (i.e., weight training, 1M steps), and an additional 0.5 hours for 1M steps of the diffusion model. However, since experiments can be run in parallel, the overall training time can be significantly reduced. The code used for training and experimentation can be found in SWG repository
Appendix E More toy examples
Figure 4: Additional toy examples with our proposed method – see Sec. 5.1 for description.

### v2

@@ B.3 Scaling of inference time in SWG
B.4 Computational resources
Our implementation is based on the jaxrl repository (jaxrl) using the JAX (jax2018github) and Flax (flax2020github) libraries. We also provide an implementation of our method in PyTorch (paszke2019pytorch). All experiments were conducted on a 12GB NVIDIA RTX 3080 Ti GPU. Training takes approximately 0.5 hours for value function learning (i.e., weight training, 1M steps), and an additional 0.5 hours for 1M steps of the diffusion model. However, since experiments can be run in parallel, the overall training time can be significantly reduced. The code used for training and experimentation can be found in the SWG repository